OpenStudy (anonymous):

Anybody good at algebra 2?

OpenStudy (anonymous):

no

OpenStudy (mathmale):

Josh-Cat, Pretty please, start by posting an actual homework problem. You'll find out soon enough whether anybody's any good at Algebra II.

OpenStudy (anonymous):

ok thank you

OpenStudy (mathmale):

Are you going to post your homework problem?

OpenStudy (anonymous):

yeah one sec

OpenStudy (anonymous):

p^2-4p-32/p+4 simply the rational expression.

OpenStudy (anonymous):

the answers are A. -p + 8; p = -4 B. p - 8; p =-4 C. -p - 8; p = -4 D. p + 8: p = -4

OpenStudy (mathmale):

Simplification of an expression does not usually involve equations, so I don't know where the equations in the possible answers came from. Are you positive you've copied the problem correctly?

OpenStudy (anonymous):

sorry and the second part says state any restriction on the expression

OpenStudy (mathmale):

I do see why p cannot equal -4: the denominator of the last term -32 ------ would be zero. But I really don't see any other restrictions. p+4 May I suggest you move on to another problem? or repost this one, double checking to ensure that you're including everything given in the original problem, and leaving more space between terms of your algebraic expression.

OpenStudy (anonymous):

p^2-4p-32 is the numerator and p+4 is the denominator.

OpenStudy (mathmale):

Oh, that changes the situation considerably. It's important you use parentheses so that it's clear what constitutes the numerator and what constitutes the denominator. your numerator is p^2-4p-32 and your denominator is p+4; that was not at all clear before. Have you tried factoring the numerator? If not, please do that now.

OpenStudy (mathmale):

Josh?

OpenStudy (mathmale):

If you factor the numerator, you'll get (p+4)(p-8). If you reduce the rational function, you'll get p-8. So? Explain in your own words what happened here, which answer is correct and why it is correct.